# A Personalised 3D+t Mesh Generative Model for Unveiling Normal Heart Dynamics

### v1

@@ Acknowledgements
This research was conducted using the UK Biobank Resource under Application Number 18545. Images were reproduced with kind permission of UK Biobank. We wish to thank all UK Biobank participants and staff. This work is supported by EPSRC DeepGeM Grant (EP/W01842X/1) and the BHF New Horizons Grant (NH/F/23/70013). K.A.M. is supported by the British Heart Foundation (FS/IPBSRF/22/27059, RE/18/4/34215) and the NIHR Imperial College Biomedical Research Centre. S.W. is supported by Shanghai Sailing Program (22YF1409300), CCF-Baidu Open Fund (CCF-BAIDU 202316) and International Science and Technology Cooperation Program under the 2023 Shanghai Action Plan for Science (23410710400). D.P.O is supported by the Medical Research Council (MC_UP_1605/13); National Institute for Health Research (NIHR) Imperial College Biomedical Research Centre; and the British Heart Foundation (RG/19/6/34387, RE/18/4/34215, CH/P/23/80008).
@@
MY.Q. and WJ.B. conceived the study. MY.Q. conducted the experiment(2-5), K.A.M. conducted the experiment(5b), MY.Q. and WJ.B. analysed the results, D.P.O. and P.M.M. provided data resources. All authors reviewed the manuscript.
Additional information

### v2

@@ Acknowledgements
This research was conducted using the UK Biobank Resource under Application Number 18545. Images were reproduced with kind permission of UK Biobank. We wish to thank all UK Biobank participants and staff. We also thank Weitong Zhang for helpful discussions on the methodology. This work is supported by EPSRC DeepGeM Grant (EP/W01842X/1) and the BHF New Horizons Grant (NH/F/23/70013). K.A.M. is supported by the British Heart Foundation (FS/IPBSRF/22/27059, RE/18/4/34215) and the NIHR Imperial College Biomedical Research Centre. S.W. is supported by Shanghai Sailing Program (22YF1409300), CCF-Baidu Open Fund (CCF-BAIDU 202316) and International Science and Technology Cooperation Program under the 2023 Shanghai Action Plan for Science (23410710400). P.M.M. acknowledges generous personal support from the Edmond J. Safra Foundation and Lily Safra, an NIHR Senior Investigator Award and the UK Dementia Research Institute, which is funded predominantly by UKRI Medical Research Council. D.P.O is supported by the Medical Research Council (MC_UP_1605/13); National Institute for Health Research (NIHR) Imperial College Biomedical Research Centre; and the British Heart Foundation (RG/19/6/34387, RE/24/130023, CH/P/23/80008).
@@
MY.Q. and WJ.B. conceived the study. MY.Q. conducted the experiments 2-5, K.A.M. conducted the experiment 5b, MY.Q. and WJ.B. analysed the results, D.P.O. and P.M.M. provided data resources. All authors reviewed the manuscript.
Competing interests